NOAA SATOPS Morning Report: Friday, December 07, 2007


Customers or users who want to call to ask about problems or report outages can use the following contacts:

- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444

- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

J341

GOES

Engineers have confirmed the cause of the attitude disturbances on GOES-12 that started after the North/South maneuver. We believe there is a leak in the oxidizer line of thruster 2B. We were successful yesterday in isolating the thruster. This evening, we are planning to go into normal attitude control mode. Engineers will be monitoring the spacecraft through out the weekend. On Monday we will determine whether to begin normal imaging and sounding operations. GOES-10 continues to support the East satellite imaging and sounding operations. Since yesterday GOES-10 GVAR and all auxiliary services, including EMWIN, were broadcast simultaneously through GOES-10 and GOES-12.

All GOES operations were nominal over the past 24 hours. No anomalies affecting product delivery or s/c health & safety. For GOES status and other information please see: http://www.oso.noaa.gov/goes/index.htm

POES

POES operations were nominal over the past 24 hours No anomalies affecting product delivery or s/c health & safety. For a detailed status and description of POES see: http://www.oso.noaa.gov/poes/index.htm

DMSP

All DMSP operations were nominal over the past 24 hours. No anomalies affecting product delivery or s/c health & safety. For more DMSP Information please see: http://dmsp.ngdc.noaa.gov/dmsp.html

Questions or comments to Ron.Mahmot@noaa.gov

Please follow SpaceRef on Twitter and Like us on Facebook.